find function f and g so that h(x)=(fg)(x)
h(x) = 6/x square +8

Answer by stanbon(48568)

find function f and g so that h(x)=(fg)(x)
h(x) = 6/x square +8
----------------------
g(x) = 1/x^2
f(x) = 6x + 8
------------------
h(x) = f[g(x)] = f[1/x^2] = 6(1/x^2) + 8 = 6/x^2 + 8
==================
Cheers,
Stan H.